# Forging GEMs: Advancing Greek NLP through Quality-Based Corpus Curation and Specialized Pre-training

### v1

@@ 3 Pretraining Corpus Composition
  • Near-deduplication across the entire corpus using the MinHashLSH [50, 26] method to eliminate semantically similar records.
For the near-deduplication stage, we employed the MinHashLSH implementation from the datasketch [^2] https://github.com/ekzhu/datasketch repository. We used 5-gram subsets, MinHash signatures comprising 128 permutations and a Jaccard similarity threshold of 0.8. These parameter choices align with those used in the Meltemi 7B pretraining pipeline [46].
Additionally, to filter out noisy sentences from the OSCAR corpus, we applied more advanced quality filtering tools such as Monocleaner [^3] https://github.com/bitextor/monocleaner. This tool combines rule-based pattern matching for detecting common anomalies with a 7-gram KenLM language model [16] trained on high-quality Greek corpora.
Dataset Size (GB) Tokens (B) Training Pairs (M) Context Language
[2] 11.00 1.68 8.65 legal Greek
Europarl [22] 0.38 0.04 0.22 legal Greek
Eurolex [7] 0.92 0.11 0.57 legal Greek
Greekparl [13] 2.90 0.56 1.87 legal Greek
Political Reports of the Supreme Court [1] 1.20 0.16 0.82 legal Greek
Raptarchis [31] 0.35 0.05 0.25 legal Greek
Wikipedia [^4] https://dumps.wikimedia.org/elwiki/ 1.30 0.17 0.79 nonlegal Greek
OSCAR [42] 41.00 5.38 24.23 nonlegal Greek–English
Pile of Law [17] 9.20 1.05 7.10 legal English
@@
4 Experimental Setup
This section presents the comprehensive experimental framework designed to pre-train and evaluate the new family of Greek Embedding Models (GEMs). The proposed methodology is structured around four key components: (i) we introduce the diverse suite of modern Transformer architectures selected for this study (Section 4.1), (ii) we describe the construction of our four distinct pre-training corpora, designed to analyze the effects of domain, data quality and bilingualism (Section 4.2), (iii) we present our systematic analysis of tokenization strategies, a critical consideration for the morphologically complex Greek language (Section 4.3), (iv) we outline the standardized training procedures and computational resources used to ensure a robust and reproducible comparison across all experiments (Section A.1).
@@ 4.1 Model Architectures
This study evaluates a comprehensive suite of modern Transformer-based encoder architectures to identify the most effective models for Greek language processing. The models were selected to represent a diverse range of innovations in the Transformer ecosystem, including highly optimized baselines (RoBERTa), efficient pre-training objectives (ELECTRA), advanced attention mechanisms (DeBERTa) and solutions for long-context processing (Longformer, ModernBERT). For each architecture, we trained a new model from scratch using default configurations, such as a Masked Language Model (MLM) probability of 15%, unless otherwise specified for a particular experiment. The core architectures investigated are:
@@
  • RoBERTa: A robustly optimized BERT pre-training approach, which improves upon the original by using a more effective training methodology, including dynamic masking and removing the Next Sentence Prediction objective [27].
@@
  • ELECTRA: An efficient architecture, which utilizes a generator-discriminator setup [9]. Instead of MLM, it employs a more sample-efficient pre-training task called Replaced Token Detection (RTD), where a discriminator model learns to identify which tokens in an input sequence were replaced by a small generator model.
@@
  • ConvBERT: A hybrid model, which enhances BERT’s efficiency by replacing some of its global self-attention heads with a novel span-based dynamic convolution [19]. This mixed-attention mechanism is designed to more efficiently model local dependencies.
@@
  • Longformer: An architecture designed to process long documents by using a sparse attention mechanism which combines local windowed attention with task-motivated global attention [6]. While the original architecture supports sequences up to 4096 tokens, we trained our version with a 1024-token context window.
@@
  • ModernBERT: A recent architecture, which modernizes BERT with several key innovations for improved efficiency and long-context handling [47]. It replaces absolute positional embeddings with Rotary Positional Embeddings (RoPE) [41], uses a GeGLU activation function [40] for better performance and employs an alternating attention strategy where layers switch between global and local sliding-window attention. While the full architecture supports sequences up to 8192 tokens, we implemented the first phase of its pre-training curriculum, focusing on a 1024-token context length.
In addition to these core architectures, we investigated several specific variations to probe the impact of key training parameters. To explore architectural scale, we trained a RoBERTa Large model to compare against its base-sized counterpart. To assess the impact of numerical precision, one RoBERTa model was trained in full precision (fp32) as a control against our standard mixed-precision (bfloat16) setup. We also experimented with the MLM probability, increasing it from the standard 15% to 20% for one RoBERTa model. This modification presents a more challenging pre-training task, potentially forcing the model to learn more robust contextual representations, which can be beneficial for complex, domain-specific greek vocabularies. Finally, for our bilingual training experiments, we utilized both the RoBERTa and XLM-RoBERTa architectures on our bilingual legal corpus.
4.2 Pre-training Corpora Configurations
From the raw data sources described in Section 3, we constructed four distinct pre-training corpora to systematically evaluate our models under different conditions. Each corpus was designed to target a specific linguistic or domain objective.
  • Legal Corpus: Our primary domain-specific corpus is composed of all Greek legal and parliamentary sources, yielding a total of approximately 16.75GB. This dataset, which comprises government gazettes, parliamentary proceedings and judicial reports, serves as the baseline for training our specialized legal language models.
  • General-Domain Corpus: To train a more broadly capable model, this corpus was constructed by combining the 17GB Legal Corpus with our cleaned general-domain sources, namely the Greek portion of Wikipedia and the OSCAR corpus. This resulted in a comprehensive dataset of 59GB, providing a rich and diverse linguistic foundation.
  • High-Quality (HQ) Repeated Corpus: To investigate the impact of data quality, an experiment was designed based on a data repetition strategy. Using an in-house data analysis tool to assess linguistic quality, a quality hierarchy was established within our legal data. The highest-quality subsets, such as the Raptarchis Legal Dictionary, were then upsampled with repetition factors of up to 4x. This factor was chosen to strike a balance between emphasizing premium legal terminology and mitigating the risk of overfitting on a narrow subset of the data, a point beyond which we observed diminishing returns in preliminary tests. This strategy expanded the original 16.75GB corpus to an effective size of 21.12GB.
  • Bilingual Legal Corpus: To facilitate cross-lingual experiments and, crucially, to investigate whether augmenting our Greek data with English text would dilute model quality, a 26GB bilingual corpus was compiled. This corpus consists of approximately 60% Greek and 40% English legal content, a ratio intentionally chosen to allocate more resources to Modern Greek, reflecting our judgment that its greater morphological complexity requires more data for effective learning. The English portion includes parallel or analogous sources to the Greek data, such as EU legislation (Eur-Lex), parliamentary proceedings (Europarl) and judicial opinions, creating a strong basis for learning cross-lingual legal representations.
@@ 4.3 Tokenization Strategy
  • Byte-Pair Encoding (BPE): An algorithm which iteratively merges the most frequent pair of adjacent tokens [39].
  • WordPiece: An algorithm which builds a vocabulary by merging pairs that maximize the likelihood of the training data [49].
  • Unigram: An algorithm which starts with a large vocabulary and progressively removes tokens to optimize a probabilistic language model [24].
To quantitatively assess the efficiency of our trained tokenizers, we selected two intrinsic metrics which directly measure the fundamental trade-off between linguistic representation and computational efficiency. These metrics were evaluated on a diverse, held-out Greek dataset of approximately 150 million tokens, spanning formal legal language, named entities and colloquialisms. The chosen metrics are: Fertility and Bytes per Token. The former is defined as the average number of subword tokens generated per word (where lower is better) while the latter measures the average number of UTF-8 bytes each token represents (where higher is better) [37]. The former indicates the degree of word fragmentation, while the latter reflects the tokenizer’s compression efficiency.
@@ 5 Numerical Experiments and Analysis
In this section, we compare the performance of the proposed class of GEM against that of the state-of-the-art Greek models: GreekBERT [23], Greek-Legal BERT [4], Greek-Legal RoBERTa [38] and Greek-Media LongFormer [52]. At this points, it is worth mentioning that the proposed class of models is composed by the architecturally advanced models: RoBERTa, ELECTRA, ConvBERT, ModernBert and LongFormer, which were training on the developed datasets: Legal, HQ Repeated and General-Domain.
In our original experiments, DeBERTa was evalued in our simulations but it reported poor performance compared to all other models; therefore, it was not included. The underperformance suggests that the DeBERTa architecture may require substantially different hyperparameter configurations or extended pretraining regimens to achieve competitive results in the Greek legal domain. A possible reason which came out from our experiments was that DeBERTa uses Unigram tokenizer, which reported poor performance. In preliminary experiments, DeBERTa was evaluated but consistently underperformed relative to other architectures and was consequently excluded from the final comparison. One potential contributing factor identified through our ablation studies is DeBERTa’s reliance on the Unigram tokenization algorithm, which demonstrated inferior performance in capturing the morphological complexity inherent to the Greek language (see Table 2).
To quantitatively assess model performance across all benchmarks, we employed three standard evaluation metrics: Accuracy (Acc), which measures the proportion of correctly classified instances across all test samples; Area Under the ROC Curve (AUC), which evaluates the model’s discriminative capacity by measuring its ability to rank positive instances higher than negative ones across varying classification thresholds; and Geometric Mean (GM), which provides a balanced performance measure that accounts for both sensitivity and specificity, proving especially valuable when dealing with class-imbalanced datasets where traditional accuracy may be misleading [34, 32].
The experimental analysis is conducted across three core natural language understanding benchmarks: Named Entity Recognition (NER), Multi-Class Legal Topic Classification (MCLTC) and Natural Language Inference (NLI). For each benchmark, we employ well-established Greek benchmark datasets, which reflect diverse linguistic and domain-specific challenges. Specifically, we utilize the Greek-Legal NER dataset for entity recognition, the Greek-Legal Code dataset for hierarchical legal topic classification and the Multi-Genre Natural Language Inference dataset for evaluating reasoning capabilities. These datasets are brief described in detail below:
  • Greek-Legal NER dataset [3]: The dataset consists of 254 issues of the Greek Government Gazette published between 2000 and 2017. It is annotated with seven entity types: legislative references, geopolitical entities, national locations, unknown locations, public locations, organizations and facilities. The annotations follow the inside–outside–beginning (IOB) format and the dataset is divided into 67.5% training, 17.5% validation and 15% test sets.
  • Greek-Legal Code dataset [31]: This dataset is based on a rich collection of Greek legislative documents published between 1834 and 2015, classified into hierarchical categories ranging from broader to more specialized legal topics. It includes Laws, Royal and Presidential Decrees, Regulations and Decisions, all retrieved from the Official Government Gazette. The dataset consists of 47 legislative volumes, each representing a main thematic topic: each volume is divided into thematic chapters, which are further subdivided into subjects. In total, the dataset contains 389 chapters and 2,285 subjects. Before splitting the dataset into training, validation and testing datasets, we removed categories (volume, chapter, subject) with fewer than 15 occurrences and ensured a balanced distribution of categories in the splits.
  • Multi-Genre Natural Language Inference dataset [48]: This dataset contains 393,000 machine-translated Greek training sentence pairs labeled as contradiction, entailment, or neutral. Each pair consists of a premise and a hypothesis and the objective is to determine whether the premise entails, contradicts, or is neutral with respect to the hypothesis. For evaluation, we used Greek development and test sets from the Cross-lingual Natural Language Inference (XNLI) corpus [10], which include 2,500 development and 5,000 test pairs, providing a reliable benchmark for assessing NLI performance in Greek.
@@
All pre-training experiments were conducted on Amazon Web Services (AWS) using either p4d.24xlarge instances (8× NVIDIA A100 GPUs, 40GB VRAM each) for standard models or p5.48xlarge instances (8× NVIDIA H100 GPUs, 80GB VRAM each) for more computationally demanding architectures such as ModernBERT and bilingual variants. To ensure fair comparison, we standardized the training regimen across all experiments using the AdamW optimizer with a linear learning rate scheduler and bfloat16 mixed precision by default, with one RoBERTa variant trained in full fp32 precision for comparative analysis. Key hyperparameters were carefully tuned for each model architecture and dataset combination, as summarized in Table 9.
@@
Tables 3, 4 and 5 report the experimental results of all transformer models on NER, MCLTC and NLI benchmarks, respectively. GEM-RoBERTa and GEM-ConvBERT exhibit the best overall performance across all three benchmarks, substantially outperforming the traditional GreekBERT, Greek-Legal BERT and Greek-Legal RoBERTa. Furthermore, GEM-ELECTRA and GEM-ModernBERT report competitive and sometimes superior performance compared to traditional Greek embedding models; therefore, validating the effectiveness of these modern architectural innovations for Greek language processing. Notably, while the Longformer variants generally exhibit the weakest performance among all evaluated architectures, the proposed GEM-Longformer consistently outperforms the Greek-Media-Longformer baseline, indicating that our improved training methodology and corpus curation strategy yield measurable gains even for architectures designed primarily for extended context processing.
@@
In addition, the analysis of the training corpus configurations reveals some interesting findings about the performance of the proposed class of models. For the NER benchmark, all GEM models achieve their best performance when trained on the HQ-Repeated corpus, with the notable exception of GEM-RoBERTa, which attains superior results using the Legal corpus. Similarly, for the MCLTC benchmark, the HQ-Repeated corpus proves most effective for all GEM variants except GEM-ConvBERT, which demonstrates the best performance with the Legal corpus. In contrast, for the NLI benchmark, all GEM models consistently achieve their best performance when trained on the General-Domain corpus, reflecting the broader linguistic diversity required for robust natural language inference capabilities. Additionally, the bilingual GEM-RoBERTa models demonstrate competitive performance across all benchmarks, occasionally achieving the best overall results. This finding is particularly significant as it indicates that the augmentation of Greek legal text with English legal data does not degrade model performance and in certain contexts, may provide complementary cross-lingual benefits that enhance the model’s representational capacity.
@@
Table 3: Performance comparison of Greek embedding models on the NER benchmark.
@@
Table 4: Performance comparison of Greek embedding models on the multi-class legal topic classification benchmark.
@@
Table 5: Performance comparison of Greek embedding models on the NLI benchmark.
For mitigating the influence of benchmarks on the evaluation process, a statistical analysis is performed for evaluating the hypothesis $H_{0}$ that all Greek embeddings models exhibit equally accurate predictions. For this purpose, non-parametric Friedman Aligned-Ranks (FAR) test [18] and the Finner post-hoc test [14] with significance level $\alpha=5\%$ are employed. The former is employed for ranking the models based on a performance metric while the latter is used for examining the existence of considerable differences and variations in their predictions, without any assumption about the distribution of the performance scores [21, 28, 33].
The statistical analysis was conducted comparing the traditional baseline models (Greek-BERT, Greek-Legal BERT, Greek-Legal RoBERTa and Greek-Media LongFormer) and the proposed GEM models in their optimal configurations. Specifically, the evaluated GEM models include: GEM-RoBERTa, GEM-ELECTRA, GEM-ConvBERT and GEM-Longformer, each trained on the General-Domain corpus and GEM-ModernBERT trained on the HQ Repeated corpus. [^5] GEM-RoBERTa, GEM-ELECTRA, GEM-ConvBERT and GEM-Longformer are available in https://huggingface.co/novelcore/models These corpus configurations were selected based on their reported performance across all evaluation benchmarks, ensuring that each GEM architecture is represented at its best capacity for a fair and comprehensive comparison against the established state-of-the-art models.

### v2

@@ 3.1 Pretraining Corpus Composition
  • Near-deduplication across the entire corpus using the MinHashLSH broder1997resemblance ; leskovec2020mining method to eliminate semantically similar records.
For the near-deduplication stage, we employed the MinHashLSH implementation from the datasketch [^2] https://github.com/ekzhu/datasketch repository. We used 5-gram subsets, MinHash signatures comprising 128 permutations and a Jaccard similarity threshold of 0.8. These parameter choices align with those used in the Meltemi 7B pretraining pipeline voukoutis2024meltemi .
Additionally, to filter out noisy sentences from the OSCAR corpus, we applied more advanced quality filtering tools such as Monocleaner [^3] https://github.com/bitextor/monocleaner. This tool combines rule-based pattern matching for detecting common anomalies with a 7-gram KenLM language model heafield2011kenlm trained on high-quality Greek corpora.
Dataset Size Tokens Training Context Language
(GB) (B) Pairs (M)
[…]
Europarl koehn2005europarl 0.38 0.04 0.22 legal Greek
Eurolex chalkidis2021multieurlex 0.92 0.11 0.57 legal Greek
Greekparl dritsa2022greek 2.90 0.56 1.87 legal Greek
Raptarchis papaloukas2021multi 0.35 0.05 0.25 legal Greek
Wikipedia [^4] https://dumps.wikimedia.org/elwiki/ 1.30 0.17 0.79 non-legal Greek
OSCAR suarez2019asynchronous 41.00 5.38 24.23 non-legal Greek–English
Pile of Law henderson2022pile 9.20 1.05 7.10 legal English
Political Reports of the Supreme Court areios2024official 1.20 0.16 0.82 legal Greek
@@
3.2 Pre-training Corpus Configurations
Based on the presented raw data sources, we constructed four distinct pre-training corpora to systematically evaluate the impact of corpus composition, domain specialization and data quality on model performance. Each corpus was strategically designed to target specific linguistic objectives and enable controlled experimentation across multiple dimensions.
  • Legal Corpus: This corpus is composed of all Greek legal and parliamentary sources, yielding a total of approximately 16.75GB. This dataset, which comprises government gazettes, parliamentary proceedings and judicial reports, serves as the baseline for training our specialized legal language models.
  • High-Quality (HQ) Repeated Corpus: To investigate the impact of data quality, an experiment was designed based on a data repetition strategy. Using an in-house data analysis tool to assess linguistic quality, a hierarchical ranking was established within our legal data. The highest-quality subsets, such as the Raptarchis Legal Dictionary, were then upsampled with repetition factors of up to 4x. This factor was chosen to strike a balance between emphasizing premium legal terminology and mitigating the risk of overfitting on a narrow subset of the data, a point beyond which we observed diminishing returns in preliminary tests. This strategy expanded the original 16.75GB corpus to an effective size of 21.12GB.
  • General-Domain Corpus: this corpus is constructed by combining the 17GB Legal Corpus with the cleaned general-domain sources, namely the Greek portion of Wikipedia and the OSCAR corpus. This resulted in a comprehensive dataset of 59GB, providing a rich and diverse linguistic foundation while maintaining legal expertise.
  • Bilingual Legal Corpus: This corpus consists of approximately 60% Greek and 40% English legal content, a ratio intentionally chosen to allocate more resources to Modern Greek, reflecting our judgment that its greater morphological complexity requires more data for effective learning. The English portion includes parallel or analogous sources to the Greek data, such as EU legislation (EUR-Lex), parliamentary proceedings (Europarl) and judicial opinions, creating a strong basis for learning cross-lingual legal representations.
Legal Corpus serves as our primary domain specific baseline, comprising exclusively Greek legal and parliamentary sources to establish a foundation for specialized legal language understanding. Motivated by the hypothesis that data quality may be more impactful than quantity for domain adaptation, we designed the HQ Repeated Corpus, which strategically upsamples premium legal resources, particularly the Raptarchis legal dictionary, balancing enhanced exposure to authoritative terminology against overfitting risks. To assess the trade-offs between domain specialization and general language proficiency, General-Domain Corpus is developed, which augments the Legal Corpus with cleaned general-domain sources from Wikipedia and OSCAR, providing broader linguistic coverage while maintaining legal expertise. Finally, to explore cross-lingual transfer and assess whether multilingual training would enhance or dilute performance for Greek legal tasks, we constructed the Bilingual Legal Corpus, combining approximately 60% Greek and 40% English legal content, with the Greek-heavy allocation reflecting our judgment that Modern Greek’s greater morphological complexity necessitates more extensive monolingual data for effective representation learning.
@@ 3.3 Model Architectures
The selection of architectures for the GEMs family was guided by three complementary principles: addressing the historical architectural stagnation in Greek NLP, systematically evaluating modern efficiency innovations, and ensuring comprehensive coverage of diverse design paradigms for encoder-based models. Along this line, we select five architectures, which collectively represent the evolution of encoder-based language models, each offering distinct advantages for processing morphologically rich languages like Greek.
@@
  • RoBERTa: A robustly optimized BERT pre-training approach, which improves upon the original by using a more effective training methodology, including dynamic masking and removing the Next Sentence Prediction objective liu2019roberta .
@@
  • ELECTRA: An efficient architecture, which utilizes a generator-discriminator setup clark2020electra . Instead of MLM, it employs a more sample-efficient pre-training task called Replaced Token Detection (RTD), where a discriminator model learns to identify which tokens in an input sequence were replaced by a small generator model.
@@
  • ConvBERT: A hybrid model, which enhances BERT’s efficiency by replacing some of its global self-attention heads with a novel span-based dynamic convolution jiang2020convbert . This mixed-attention mechanism is designed to more efficiently model local dependencies.
@@
  • Longformer: An architecture designed to process long documents by using a sparse attention mechanism which combines local windowed attention with task-motivated global attention beltagy2020longformer . While the original architecture supports sequences up to 4096 tokens, we trained our version with a 1024-token context window.
@@
  • ModernBERT: A recent architecture, which modernizes BERT with several key innovations for improved efficiency and long-context handling warner2024modernbert . It replaces absolute positional embeddings with Rotary Positional Embeddings (RoPE) su2021roformer , uses a GeGLU activation function shazeer2020glu for better performance and employs an alternating attention strategy where layers switch between global and local sliding-window attention. While the full architecture supports sequences up to 8192 tokens, we implemented the first phase of its pre-training curriculum, focusing on a 1024-token context length.
@@ 3.4 Tokenization Strategy
  • Byte-Pair Encoding (BPE): An algorithm which iteratively merges the most frequent pair of adjacent tokens sennrich2016neural .
  • WordPiece: An algorithm which builds a vocabulary by merging pairs that maximize the likelihood of the training data wu2016google .
  • Unigram: An algorithm which starts with a large vocabulary and progressively removes tokens to optimize a probabilistic language model kudo2018subword .
To quantitatively assess the efficiency of our trained tokenizers, we selected two intrinsic metrics which directly measure the fundamental trade-off between linguistic representation and computational efficiency. These metrics were evaluated on a diverse, held-out Greek dataset of approximately 150 million tokens, spanning formal legal language, named entities and colloquialisms. The chosen metrics are: Fertility and Bytes per Token. The former is defined as the average number of subword tokens generated per word (where lower is better) while the latter measures the average number of UTF-8 bytes each token represents (where higher is better) rust2021good . The former indicates the degree of word fragmentation, while the latter reflects the tokenizer’s compression efficiency.
@@ 4 Numerical Experiments and Analysis
In this section, we compare the performance of the proposed family of GEMs against that of the state-of-the-art Greek models: GreekBERT koutsikakis2020greekbert , Greek-Legal BERT apostolopoulou2022nlp , Greek-Legal RoBERTa saketos2024greeklegalroberta and Greek-Media LongFormer 10288436 . At this point, it is worth mentioning that the proposed class of models consists of the architecturally advanced RoBERTa, ELECTRA, ConvBERT, ModernBert and LongFormer as well as the well stablished RoBERTa and LongFormer, which were trained on the developed datasets: Legal, HQ Repeated and General-Domain. For all pre-training experiments, the data was split into 90% training and 10% validation sets, with early stopping employed to prevent overfitting and ensure optimal model generalization.
@@
All pre-training experiments were conducted on Amazon Web Services (AWS) using either p4d.24xlarge instances (8× NVIDIA A100 GPUs, 40GB VRAM each) for standard models or p5.48xlarge instances (8× NVIDIA H100 GPUs, 80GB VRAM each) for more computationally demanding architectures such as ModernBERT and bilingual variants. The implementation was developed in Python using the PyTorch deep learning framework and the Hugging Face Transformers library. To ensure fair comparison, we standardized the training regimen across all experiments using the AdamW optimizer with a linear learning rate scheduler and bfloat16 mixed precision by default, with one RoBERTa variant trained in full fp32 precision for comparative analysis. Key hyperparameters were carefully tuned for each model architecture and dataset combination. Detailed information about the complete hyperparameter configurations, training time and computational requirements for all experiments can be found in Appendix A.
In our original experiments, DeBERTa was evaluated in our simulations but it reported poor performance compared to all other models; hence, it was not included. The underperformance suggests that the DeBERTa architecture may require substantially different hyperparameter configurations or extended pretraining regimens to achieve competitive results in the Greek legal domain. One potential contributing factor identified through our ablation studies was that DeBERTa uses Unigram tokenizer, which reported considerably poor performance in capturing the morphological complexity inherent to the Greek language (see Appendix A).
In the sequel, we evaluate the performance of all pre-trained models across three core natural language understanding benchmarks: Named Entity Recognition (NER), Multi-Class Legal Topic Classification (MCLTC) and Natural Language Inference (NLI). For each benchmark, we employ well-established Greek benchmark datasets, which reflect diverse linguistic and domain-specific challenges. Specifically, we utilize the Greek-Legal NER dataset for entity recognition, the Greek-Legal Code dataset for hierarchical legal topic classification and the Multi-Genre Natural Language Inference dataset for evaluating reasoning capabilities. These datasets are brief described in detail below:
  • Greek-Legal NER dataset Angelidis2018NamedER : The dataset consists of 254 issues of the Greek Government Gazette published between 2000 and 2017. It is annotated with seven entity types: legislative references, geopolitical entities, national locations, unknown locations, public locations, organizations and facilities. The annotations follow the inside–outside–beginning (IOB) format and the dataset is divided into 67.5% training, 17.5% validation and 15% test sets.
  • Greek-Legal Code dataset papaloukas2021multi : This dataset is based on a rich collection of Greek legislative documents published between 1834 and 2015, classified into hierarchical categories ranging from broader to more specialized legal topics. It includes Laws, Royal and Presidential Decrees, Regulations and Decisions, all retrieved from the Official Government Gazette. The dataset consists of 47 legislative volumes, each representing a main thematic topic: each volume is divided into thematic chapters, which are further subdivided into subjects. In total, the dataset contains 389 chapters and 2,285 subjects. Before splitting the dataset into training, validation and testing datasets, we removed categories (volume, chapter, subject) with fewer than 15 occurrences and ensured a balanced distribution of categories in the splits.
  • Multi-Genre Natural Language Inference dataset williams2017broad : This dataset contains 393,000 machine-translated Greek training sentence pairs labeled as contradiction, entailment, or neutral. Each pair consists of a premise and a hypothesis and the objective is to determine whether the premise entails, contradicts, or is neutral with respect to the hypothesis. For evaluation, we used Greek development and test sets from the Cross-lingual Natural Language Inference (XNLI) corpus conneau2018xnli , which include 2,500 development and 5,000 test pairs, providing a reliable benchmark for assessing NLI performance in Greek.
To quantitatively assess model performance across all benchmarks, we employed three standard evaluation metrics: Accuracy (Acc), which measures the proportion of correctly classified instances across all test samples; Area Under the ROC Curve (AUC), which evaluates the model’s discriminative capacity by measuring its ability to rank positive instances higher than negative ones across varying classification thresholds; and Geometric Mean (GM), which provides a balanced performance measure that accounts for both sensitivity and specificity, proving especially valuable when dealing with class-imbalanced datasets where traditional accuracy may be misleading pintelas2024mobilenet ; pintelas2025textnex .
@@
Tables 3, 4 and 5 report the experimental results of all transformer models on NER, MCLTC and NLI benchmarks, respectively. It is worth mentioning that the first two columns refer to the model’s name and the utilized corpus to pre-train the model. GEM-RoBERTa and GEM-ConvBERT exhibit the best overall performance across all three benchmarks, substantially outperforming the traditional GreekBERT, Greek-Legal BERT and Greek-Legal RoBERTa. Furthermore, GEM-ELECTRA and GEM-ModernBERT report competitive and sometimes superior performance compared to traditional Greek embedding models; therefore, validating the effectiveness of these modern architectural innovations for Greek language processing. Notably, while the Longformer variants generally exhibit the weakest performance among all evaluated architectures, the proposed GEM-Longformer consistently outperforms the Greek-Media-Longformer baseline, indicating that our improved training methodology and corpus curation strategy yield measurable gains even for architectures designed primarily for extended context processing.
@@
In addition, the analysis of the training corpus configurations reveals some interesting findings about the performance of the proposed class of models. For the NER benchmark, all GEMs models achieve their best performance when trained on the HQ-Repeated corpus, with the notable exception of GEM-RoBERTa, which attains superior results using the Legal corpus. Similarly, for the MCLTC benchmark, the HQ-Repeated corpus proves most effective for all GEM variants except GEM-ConvBERT, which demonstrates the best performance with the Legal corpus. In contrast, for the NLI benchmark, all GEM models consistently achieve their best performance when trained on the General-Domain corpus, reflecting the broader linguistic diversity required for robust natural language inference capabilities. Additionally, both bilingual models, GEM-RoBERTa and GEM-XLM-RoBERTa demonstrate competitive performance across all benchmarks, occasionally achieving the best overall results, with GEM-RoBERTa present slightly better performance.
@@
Table 3: Performance comparison of Greek embedding models on the NER benchmark.Underline indicates the best performance for each GEM variant, while bold indicates the best overall performance for each metric
@@
Table 4: Performance comparison of Greek embedding models on the multi-class legal topic classification benchmark. Underline indicates the best performance for each GEM variant, while bold indicates the best overall performance for each metric
@@
Table 5: Performance comparison of Greek embedding models on the NLI benchmark. Underline indicates the best performance for each GEM variant, while bold indicates the best overall performance for each metric
Summarizing, the numerical experiments reveal the following noteworthy findings regarding corpus composition strategies. Firstly, the HQ Repeated corpus demonstrates consistent effectiveness, emerging as the (near-)optimal configuration for most GEM architectures across the majority of tasks. This superior performance is attributed to two complementary factors: the corpus comprises predominantly formal legal texts from official sources, which exhibit greater morphological richness and linguistic precision compared to web-crawled general-domain data; and the strategic upsampling of premium legal resources, particularly the Raptarchis Legal Dictionary, which provides authoritative definitions and canonical usage patterns of specialized legal terminology. Through controlled repetition of such high-quality, domain-specific content, more robust representations of legal concepts are developed without the noise inherent in web-crawled corpora. This data-centric approach demonstrates that for morphologically rich languages, training data quality and domain-relevance significantly outweigh raw corpus size, a finding with practical implications for resource, constrained language modeling. Secondly, the competitive performance of bilingual GEM-RoBERTa models demonstrates that augmenting Greek legal text with English legal content does not degrade model quality for Greek-specific tasks. In several cases, superior performance is achieved compared to monolingual variants, indicating that cross-lingual exposure to parallel legal concepts provides complementary representational benefits. This finding challenges the assumption that strict monolingual training is always optimal for moderately-resourced languages and suggests that high-quality English resources can effectively enhance specialized domain capabilities.
For mitigating the influence of benchmarks on the evaluation process, a statistical analysis is performed for evaluating the hypothesis $H_{0}$ that all Greek embeddings models exhibit equally accurate predictions. For this purpose, non-parametric Friedman Aligned-Ranks (FAR) test hodges2011rank and the Finner post-hoc test finner1993monotonicity with significance level $\alpha=5\%$ are employed. The former is employed for ranking the models based on a performance metric while the latter is used for examining the existence of considerable differences and variations in their predictions, without any assumption about the distribution of the performance scores kiriakidou2024mutual ; livieris2024c ; pintelas2024adaptive .
The statistical analysis was conducted comparing the traditional baseline models (Greek-BERT, Greek-Legal BERT, Greek-Legal RoBERTa and Greek-Media LongFormer) and the proposed GEMs in their configurations, which reported the best overall performance. Specifically, the evaluated GEM models include: GEM-RoBERTa, GEM-ELECTRA, GEM-ConvBERT and GEM-Longformer, each trained on the General-Domain corpus and GEM-ModernBERT trained on the HQ Repeated corpus. [^5] GEM-RoBERTa, GEM-ELECTRA, GEM-ConvBERT and GEM-Longformer are available in https://huggingface.co/novelcore/models These corpus configurations were selected based on their reported performance across all evaluation benchmarks, ensuring that each model is represented at its best capacity for a fair and comprehensive comparison against the established state-of-the-art models.